1. Let a and b be elements of a group G . Prove that ab and ba have the same order.
2. Show by example that the product of elements of nite order in a group need not
have nite order. What if the group is abelian?
Answer #1

element Solution : be any eleme Let a and b group G. of a To show ocab) = olbal repe where o cab) onden auf ab we know that o& Let A-(07 EGLCR) B- [! ] E GLCR) AP: A:7[:] = [0] = 1 = O(A)= 2 . Also Oz B.6=[ J[:] - [6]1 = o(B)=2 B²- l = O (A)=2 o(B) =for any new This (AB) - AB is I not of finite order 26: If G is a belian group. Then order of the product of the two finite o